OpenStudy (anonymous):

Write the quadratic equation in factored form. Be sure to write the entire equation. x2 + x - 12 = 0

OpenStudy (anonymous):

notice that x = 4x -3x use that fact to factor the equation

OpenStudy (anonymous):

\[-b \pm \sqrt{b^2-4ac}\]----------------\[2a\]\[-1\pm \sqrt{1^2-4*1*-12}\]---------------------\[2*1\]\[-1\pm \sqrt{1+12}\]--------------\[2\]

OpenStudy (anonymous):

@mathfrick 4*12 = 48 you made a slight mistake there.

OpenStudy (anonymous):

try factoring as \[(x+4)(x-3)=0\]

OpenStudy (anonymous):

\[-b \pm \sqrt{1+48}\]------\[2\]

OpenStudy (haseeb96):